OpenStudy (anonymous):

A.Write the expression 243^-2/5 in radical expression. B.Evaluate the radical expression

OpenStudy (anonymous):

your exponent is \(-\frac{2}{5}\) which has three parts a numerator of 2 which is the power, a denominator of 5 with means the fifth root and a minus sign which means take the reciprocal

OpenStudy (anonymous):

since \(243=3^5\) the fifth root of 435 is \(3\) then \(3^2=9\) and the reciprocal of 9 is \(\frac{1}{9}\)

OpenStudy (anonymous):

so the answer would be 243=\[3^{5}\] and \[\frac{ 1 }{ 9 }\]